6-29. Upon hearing that Ross White (see Problems 6-27 and 6-28) is considering producing the brackets in-house, the vendor has notified Ross that the purchase price would drop from $15 per bracket to $14.50 per bracket if Ross would purchase the brackets in lots of 1,000. Lead times, however, would increase to 3 days for this larger quantity. A) What is the total annual inventory cost plus purchase cost if Ross buys the brackets in lots of 1,000 at $14.50 each? B) If Ross does buy in lots of 1,000 brackets, what is the new ROP? C) Given the options of purchasing the brackets at $15 each, producing them in-house at $14.80, and taking advantage of the discount, what is your recommendation to Ross White?

Answer in EXCEL
Answer #1

Please give positive rating your feedback is valuable to me.

In case , any problem please leave a comment.

Answer

A. Average Inventory = 1000/2 = 500 Average Inventory holding Cost = 500 * 0.1 * 14.5 = 725 Annual Ordering Cost = (2500 * 18(B) New ROP = Lead Time * Demand (3/365) * 2500 = 20.5479 units =C) Total inventory cost for Ross White in each of the 3 cases are: Purchasing the brackets at $15 each = (250*1.5) / 2 + (250